# Data curation, Methodology, Writing - original draft

### v3

@@ 5.3 Hyperparameter Analysis
Neighborhood Slice Number. Since the proposed MOSformer is a 2.5D-based model, it requires neighborhood slices as additional inputs, as illustrated in Section 3. Thus, the number of neighborhood slices ($s$) is an important hyperparameter. Table 6 reports quantitative results for three different $s$ parameters. It can be observed that segmentation performance initially increases and then decreases with an increasing value of $s$. Evidently, information from inter-slice enables our model to perceive partial structures of 3D medical volumes. However, a peculiar phenomenon emerges: segmentation performance of the model with $s=2$ is worse than that with $s=1$. Similar observations have been reported in [18]. One possible explanation is that the most valuable inter-slice information is derived from adjacent slices. Introducing non-adjacent slices may bring redundant information, which contributes negatively to model performance. Additionally, as $s$ increases, the computational costs of our model also escalate. Based on the above observations, $s=1$ is the most practical choice for our model.
@@
2D UNet [12] [MICCAI’15] 17.26 30.74 0.67
TransUNet [23] [MedIA’24] 93.23 24.73 5.69
MISSformer [43] [TMI’23] 35.45 7.28 7.20
3D UNETR [44] [WACV’22] 92.62 82.63 5.39
nnFormer [46] [TIP’23] 149.13 246.10 10.13
2.5D CAT-Net [28] [TMI’22] 220.16 121.83 21.34
@@
Multi-scale Inter-slice Fusion. Multi-scale learning enables deep models to capture global spatial information and local contextual details. This conclusion has been supported by many studies [23][22][42]. In this paper, we further investigate multi-scale learning by incorporating inter-slice fusion. Table 7 presents results derived from four different inter-slice fusion configurations. Our default model achieves significant performance improvements, such as $1.11\%\sim+2.63\%$ gains in DSC on the multiorgan segmentation (Synapse) dataset and $0.25\%\sim+0.56\%$ gains in DSC on the automatic cardiac diagnosis (ACDC) dataset. With more scales of inter-slice information fused, MOSformer demonstrates an enhanced ability to comprehend global shapes and anatomical details within segmentation targets. This enhancement facilitates precise localization of semantic regions, resulting in higher DSC, and accurate classification of category boundaries, reflected in smaller HD95.
@@ 5.4 Model Complexity
Table 8 presents a comparison of five medical image segmentation models with MOSformer across various dimensions, including model parameters, floating-point operations per second (FLOPs), and the average time required for segmenting individual cases. MOSformer maintains a relatively small size (77.09 M) compared with 3D-based and 2.5D-based models. Furthermore, MOSformer exhibits an inference speed only half that of nnFormer [46], even surpassing 2D-based TransUNet [23] and MISSformer [43]. These results indicate MOSformer can achieve a favorable trade-off between model complexity and segmentation performance.

### v4

@@ 5.3 Hyperparameter Analysis
Neighborhood Slice Number. Since the proposed MOSformer is a 2.5D-based model, it requires neighborhood slices as additional inputs, as illustrated in Section 3. Thus, the number of neighborhood slices ($s$) is an important hyperparameter. Table 6 reports quantitative results for three different $s$ parameters. It can be observed that segmentation performance initially increases and then decreases with an increasing value of $s$. Evidently, information from inter-slice enables our model to perceive partial structures of 3D medical volumes. However, a peculiar phenomenon emerges: segmentation performance of the model with $s=2$ is worse than that with $s=1$. Similar observations have been reported in [zhang2022bridging]. One possible explanation is that the most valuable inter-slice information is derived from adjacent slices. Introducing non-adjacent slices may bring redundant information, which contributes negatively to model performance. Additionally, as $s$ increases, the computational costs of our model also escalate. Based on the above observations, $s=1$ is the most practical choice for our model.
@@
2D UNet [ronneberger2015u] [MICCAI’15] 17.26 30.74 0.67
TransUNet [chen2024transunet] [MedIA’24] 93.23 24.73 5.69
MISSformer [huang2023missformer] [TMI’23] 35.45 7.28 7.20
3D UNETR [hatamizadeh2022unetr] [WACV’22] 92.62 82.63 5.39
nnFormer [zhou2023nnformer] [TIP’23] 149.13 246.10 10.13
2.5D CAT-Net [hung2022cat] [TMI’22] 220.16 121.83 21.34
@@
Multi-scale Inter-slice Fusion. Multi-scale learning enables deep models to capture global spatial information and local contextual details. This conclusion has been supported by many studies [chen2024transunet][cao2022swin][you2022class]. In this paper, we further investigate multi-scale learning by incorporating inter-slice fusion. Table 7 presents results derived from four different inter-slice fusion configurations. Our default model achieves significant performance improvements, such as $1.11\%\sim+2.63\%$ gains in DSC on the multiorgan segmentation (Synapse) dataset and $0.25\%\sim+0.56\%$ gains in DSC on the automatic cardiac diagnosis (ACDC) dataset. With more scales of inter-slice information fused, MOSformer demonstrates an enhanced ability to comprehend global shapes and anatomical details within segmentation targets. This enhancement facilitates precise localization of semantic regions, resulting in higher DSC, and accurate classification of category boundaries, reflected in smaller HD95.
@@ 5.4 Model Complexity
Table 8 presents a comparison of five medical image segmentation models with MOSformer across various dimensions, including model parameters, floating-point operations per second (FLOPs), and the average time required for segmenting individual cases. MOSformer maintains a relatively small size (77.09 M) compared with 3D-based and 2.5D-based models. Furthermore, MOSformer exhibits an inference speed only half that of nnFormer [zhou2023nnformer], even surpassing 2D-based TransUNet [chen2024transunet] and MISSformer [huang2023missformer]. These results indicate MOSformer can achieve a favorable trade-off between model complexity and segmentation performance.